Given are the following equations for lentils:

Qs = 2P and Qd=300-P

a. Government imposes a price ceiling of Rs.50 on lentils. Calculate the consumer and producer 

surplus at the ceiled price?

b. Calculate the dead weight loss. 


1
Expert's answer
2022-01-17T12:27:08-0500

Qs = 2P or Ps = 0.5Q, Qd = 300 - P or Pd = 300 - Q.

In equilibrium Qd = Qs, so:

300 - P = 2P,

P = 100,

Q = 2×100 = 200 units.

a. If government imposes a price ceiling of Rs.50 on lentils, then the consumer and producer surplus at the ceiled price are:

Q = 2×50 = 100 units.

CS = 100×((300 - 50) + (250 - 50))/2 = 22,500.

PS = 0.5×100×50 = 2,500.

b. The dead weight loss is:

DWL = 0.5×(200 - 100)×(250 - 50) = 10,000.